# On the microscopics of proximity effects in one-dimensional superconducting hybrid systems

### v1

@@ II The NEGF Model
Figure 1: Pairing amplitude profile across the SN junction in a hybrid nanowire setup. (a) Schematic of the device showing the effective one-dimensional nanowire with the SN junction at $x=0$. (b) One-dimensional tight-binding model of the nanowire within the NEGF framework, showing the lattice discretization $a$, hopping strength $t_{0}$, contact self energies $\Sigma_{S/N}$, and the number of lattice sites considered in the Hamiltonian $L_{s/n}$. (c) Pair amplitude across the SN junction exhibiting the non-trivial decay of superconducting correlations inside the normal region. $\Delta=1$meV and $\mu=10$meV. The pair amplitude at the interface $F_{0}\equiv F(x=0)$, and the effective coherence length $\xi_{SN}$ for the SN junction are indicated on the plot.
@@
Figure 2: In and out of the Andreev approximation. Top row (a-b): The chemical potential is varied in the range $\mu/\Delta\in\{1.2,3,5,10,15\}$ with a constant superconducting order $\Delta=1$meV. (a) Pair amplitude computed across the SN junction for different values of $\mu/\Delta$. (b) Effective coherence length plotted as a function of $\mu/\Delta$ with a fitted curve of $\xi_{SN}(\mu/\Delta)=2.7\log{\mu/\Delta}+29.4$(nm) superposed. Bottom row (c-d): The superconducting order parameter is varied in the range of $\Delta\in\{0.1,0.5,1,2,3,4,5\}$meV for a constant chemical potential $\mu=10$meV. (d) Pair amplitude computed across the SN junction for different values of $\Delta$. (e) Effective coherence length as a function of $\Delta$ with a fitted curve of $15.7/\Delta^{2}+17.2/\Delta-0.94~{}$(nm).
@@
Figure 5: Disorder induced crossover (a) Schematic of the disorder profile in the normal region, with $\sigma(x)$ denoting the variance of the zero-mean uncorrelated potential at each location $x\geq 0$, and the spread coefficient $\sigma_{D}$ denoting the spatial extent of $\sigma(x)$ into the normal region (b) A sketch of the induced correlations profile under disorder as a function of $\gamma$ (decay coefficient specifying the order of the potential $V/t_{0}\sim 10^{-\gamma}$) and $\sigma_{D}$ (spread of variance profile) (c) Pair amplitude for (i) $\sigma_{D}=5,~{}\gamma\in\{3,1\}$ (ii) $\sigma_{D}=25,~{}\gamma\in\{3,1\}$, (iii) $\sigma_{D}=50,~{}\gamma\in\{3,1.75\}$ and (iv) $\sigma_{D}=100,~{}\gamma\in\{3,2\}$ with the red curves shown in the inset with polynomial/exponential fits to the tails and corresponding to the marked points on the schematic. The vertical dotted line denotes the variance of the disorder profile $a\sigma_{D}$.
@@ III.1 Induced superconducting correlations in SN junctions
We now discuss the behavior of the pair amplitude evaluated using the NEGF method discussed in the previous section for the prototypical SN junction of Fig. 1(a). The result is shown in Fig. 1(c) for the case where the chemical potential is $\mu=10$meV and the order parameter is $\Delta=1$meV, thereby satisfying the Andreev approximation of $\mu/\Delta\gg 1$. We consider the system at zero temperature, and with no voltage/thermal drive on the baths. Deep in the superconducting side $x\to-\infty$, the pair amplitude attains a constant non-zero value signifying the bulk correlations in the superconductor. As we move closer to the interface $x\to 0$, the amplitude shows a decaying profile into the normal region, and decays to zero deep in the normal region $x\to\infty$. We denote the value of $F(x)$ at the junction as $F_{0}\equiv F(x=0)$. The decay of the pair amplitude into the normal region is found to be power law of the exact form,
@@
where $\xi_{SN}$ is thus interpreted as an effective coherence length of the junction, defining a length-scale for algebraically decaying correlations ($F(x)\to F_{0}/2\text{ as }x\to\xi_{SN}$). The power-law decay of the pair amplitude, in turn, indicates a long-range penetration of the induced correlations into the normal region. We attribute this to the absence of disorder in the normal region, which leads to an infinite coherence length, since the retroreflected particle does not lose the phase information regarding the superconductor by scattering events in the non-ordered region. This finding is in agreement with the power-law order parameter decay found in the case of clean nanowires at zero temperature in Ref. [63] where a self-consistent Bogoliubov-de Gennes (BdG) approach is employed. In doing so, this establishes the correctness of our method thereby benchmarking with traditional self-consistent calculations.
@@ III.2 Effect of the Andreev approximation
The results for the pair amplitude where we tune in and out of the Andreev approximation are shown in the top row of Fig. 2. We see that, when for a fixed $\Delta=1$meV as we increase the ratio $\mu/\Delta$ in the set $\{1.2,3,5,10,25\}$, the bulk value of pair amplitude inside the superconductor decreases, as the in-gap excitations are decreased along with an increase in the chemical potential. This further implies that the induced amplitude inside the normal region decreases upon tuning into the Andreev approximation. But, upon careful inspection, it can be shown that the effective coherence length of the junction $\xi_{SN}$ as defined in Sec. III.1, increases with the value of $\mu/\Delta$. This is shown in Fig. 2(b), where the growth of the effective coherence length is further shown to be logarithmic in the parameter $\mu/\Delta$. This implies that the pair correlations leaking into the normal region increase as we go deeper into the Andreev approximation regime, which is in agreement with the sub-gap conductance picture.
@@
Finally, we outline how the effective coherence length of the junction evolves as the superconducting order is increased. This is shown in the bottom row of Fig. 2. We plot the induced pair amplitude for a fixed chemical potential of $\mu=10$meV for the different order parameter values $\Delta\in\{0.1,0.5,1,2,3,4,5\}$meV in Fig. 2(c). We see an increasing value of the pair amplitude in the bulk superconductor and the induced amplitude in the normal region as $\Delta$ is increased. The effective coherence lengths are then plotted in Fig. 2(d). We note an algebraic decay in $1/\Delta$ of the effective coherence length. This highlights an important consideration when designing hybrid devices that leverage the proximity effect: two different values of the nearby superconducting order, $\Delta_{1}$ and $\Delta_{2}$, can yield nearly identical sub-gap conductance in the near-pristine limit (e.g., when $\mu/\Delta_{1},\mu/\Delta_{2}\gg 1$, as in the case of $\Delta_{1}=0.1$ meV and $\Delta_{2}=0.5$ meV here). However, they will induce superconducting order that propagates differently within the normal region, as reflected in the effective coherence length. Moreover, although this algebraic dependence resembles the coherence length of a BCS superconductor ($\xi_{S}\sim 1/\Delta$), it must be noted that the two quantities describe fundamentally different phenomena. The BCS coherence length represents the average size of a Cooper pair within the superconductor, whereas the effective coherence length of the junction here quantifies the spatial extent of the induced correlations within the normal region.
@@ III.4 Disorder in the SN junction
A schematic diagram of the correlation patterns in the system as a function of the decay rate $\gamma$ and the variance spread parameter $\sigma_{D}$ is shown in Fig. 5(b). The previous results in Sec. III.1 were in the $\gamma\to\infty$ and $\sigma_{D}\to 0$ limit where we observed a power-law, or more generally algebraic decay in $1/r$ of the pair amplitude as a function of $r$. In the other limit of $\gamma\to 0$ and $\sigma_{D}\to\infty$, we expect the coherence properties of the normal region to be destroyed and thus infer a fully suppressed proximity-effect. We can indeed show, with limited computational resources on a finite size system, the transition between these two limits, as indicated in Fig. 5(b). In a more rigorous sense, we expect this to be a phase transition in the steady state $\rho_{\textsc{SS}}$ of the system, which we probe through the pair correlations in $\rho_{\textsc{SS}}$ using the methods developed before.

### v2

@@ II The NEGF Model
Figure 1: Pairing amplitude profile across the SN junction in a hybrid nanowire setup. (a) Schematic of the device showing the effective one-dimensional nanowire with the SN junction at $x=0$. (b) One-dimensional tight-binding model of the nanowire within the NEGF framework, showing the lattice discretization $a$, hopping strength $t_{0}$, contact self energies $\Sigma_{S/N}$, and the number of lattice sites considered in the Hamiltonian $L_{s/n}$. (c) Pair amplitude across the SN junction exhibiting the non-trivial decay of superconducting correlations inside the normal region. $\Delta=1$meV and $\mu=10$meV. The pair amplitude at the interface $F_{0}\equiv F(x=0)$, and the effective coherence length $\xi_{SN}=44(1)$nm for the SN junction are indicated on the plot.
@@
Figure 2: In and out of the Andreev approximation. Top row (a-b): The chemical potential is varied in the range $\mu/\Delta\in\{1.2,3,5,10,15\}$ with a constant superconducting order $\Delta=1$meV. (a) Pair amplitude computed across the SN junction for different values of $\mu/\Delta$. (b) Effective coherence length plotted as a function of $\mu/\Delta$ with a fitted curve of $\xi_{SN}(\mu/\Delta)=6.8\sqrt{\mu/\Delta}+28.7$(nm) superposed. Bottom row (c-d): The superconducting order parameter is varied in the range of $\Delta\in\{0.1,0.5,1,2,3,4,5\}$meV for a constant chemical potential $\mu=10$meV. (d) Pair amplitude computed across the SN junction for different values of $\Delta$. (e) Effective coherence length as a function of $\Delta$ with a fitted curve of $15.1+26.2/\Delta-1.5/\Delta^{2}~{}$(nm).
@@
Figure 5: Disorder induced crossover (a) Schematic of the disorder profile in the normal region, with $\sigma(x)$ denoting the variance of the zero-mean uncorrelated potential at each location $x\geq 0$, and the spread coefficient $\sigma_{D}$ denoting the spatial extent of $\sigma(x)$ into the normal region (b) A map of the induced correlations profile under disorder as a function of $\gamma$ (decay coefficient specifying the order of the potential $V/t_{0}\sim 10^{-\gamma}$) and $\sigma_{D}$ (spread of variance profile) consisting of four data points and a sketch of the crossover line separating the two decay behavior (c) Pair amplitude for (i) $\sigma_{D}=5,~{}\gamma\in\{3,1\}$ (ii) $\sigma_{D}=25,~{}\gamma\in\{3,1\}$, (iii) $\sigma_{D}=50,~{}\gamma\in\{3,1.75\}$ and (iv) $\sigma_{D}=100,~{}\gamma\in\{3,2\}$ with the red curves shown in the inset with polynomial/exponential fits to the tails and corresponding to the marked points on the schematic. The vertical dotted line denotes the variance of the disorder profile $a\sigma_{D}$.
@@ III.1 Induced superconducting correlations in SN junctions
We now discuss the behavior of the pair amplitude evaluated using the NEGF method discussed in the previous section for the prototypical SN junction of Fig. 1(a). The result is shown in Fig. 1(c) for the case where the chemical potential is $\mu=10$meV and the order parameter is $\Delta=1$meV, thereby satisfying the Andreev approximation of $\mu/\Delta\gg 1$. We consider the system at zero temperature, and with no voltage/thermal drive on the baths. Deep in the superconducting side $x\to-\infty$, the pair amplitude attains a constant non-zero value signifying the bulk correlations in the superconductor. As we move closer to the interface $x\to 0$, the amplitude shows a decaying profile into the normal region, and decays to zero deep in the normal region $x\to\infty$. We denote the value of $F(x)$ at the junction as $F_{0}\equiv F(x=0)$. The decay of the pair amplitude into the normal region is found to be power law of the form,
@@
where $\xi_{SN}$ is thus interpreted as an effective coherence length of the junction, defining a length-scale for algebraically decaying correlations ($F(x)\to F_{0}/2\text{ as }x\to\xi_{SN}$). In Fig. 1 we find $\xi_{SN}=44(1)$nm. Moving further, the effective junction coherence length is obtained by fitting to the form of Eq. 11 unless otherwise specified. The power-law decay of the pair amplitude, in turn, indicates a long-range penetration of the induced correlations into the normal region. We attribute this to the absence of disorder in the normal region, which leads to an infinite coherence length, since the retroreflected particle does not lose the phase information regarding the superconductor by scattering events in the non-ordered region. This finding is in agreement with the power-law order parameter decay found in the case of clean nanowires at zero temperature in Ref. [63] where a self-consistent Bogoliubov-de Gennes (BdG) approach is employed. In doing so, this establishes the correctness of our method thereby benchmarking with traditional self-consistent calculations.
@@ III.2 Effect of the Andreev approximation
The results for the pair amplitude where we tune in and out of the Andreev approximation are shown in the top row of Fig. 2. We see that, when for a fixed $\Delta=1$meV as we increase the ratio $\mu/\Delta$ in the set $\{1.2,3,5,10,25\}$, the bulk value of pair amplitude inside the superconductor decreases, as the in-gap excitations are decreased along with an increase in the chemical potential. This further implies that the induced amplitude inside the normal region decreases upon tuning into the Andreev approximation. But, upon careful inspection, it can be shown that the effective coherence length of the junction $\xi_{SN}$ increases with the value of $\mu/\Delta$. This is shown in Fig. 2(b), where the growth of the effective coherence length is $\sim\sqrt{\mu/\Delta}$. This is extracted by fitting $\sim F_{0}\xi_{SN}/x$ to the tail of the pair amplitude curves at $x\gtrsim 200$nm, as we expect the coherence length away from the interface to be related to the only length-scale present in the ballistic system—the Fermi wavelength, which depends on $\sqrt{\mu}$—upto finite size effects. The numerical coefficients are phenomenological, and expected to change with the microscopic parameters. Moreover, this increase in coherence implies that the pair correlations leaking into the normal region increase as we go deeper into the Andreev approximation regime, which is in agreement with the sub-gap conductance picture.
@@
Finally, we outline how the effective coherence length of the junction evolves as the superconducting order is increased. This is shown in the bottom row of Fig. 2. We plot the induced pair amplitude for a fixed chemical potential of $\mu=10$meV for the different order parameter values $\Delta\in\{0.1,0.5,1,2,3,4,5\}$meV in Fig. 2(c). We see an increasing value of the pair amplitude in the bulk superconductor and the induced amplitude in the normal region as $\Delta$ is increased. The effective coherence lengths are then plotted in Fig. 2(d). We note an algebraic decay in $1/\Delta$ of the effective coherence length, which goes as $\sim 1/\Delta$ for large $\Delta$. This highlights an important consideration when designing hybrid devices that leverage the proximity effect: two different values of the nearby superconducting order, $\Delta_{1}$ and $\Delta_{2}$, can yield nearly identical sub-gap conductance in the near-pristine limit (e.g., when $\mu/\Delta_{1},\mu/\Delta_{2}\gg 1$, as in the case of $\Delta_{1}=0.1$ meV and $\Delta_{2}=0.5$ meV here). However, they will induce superconducting order that propagates differently within the normal region, as reflected in the effective coherence length. Moreover, although this algebraic dependence resembles the coherence length of a BCS superconductor ($\xi_{S}\sim 1/\Delta$), it must be noted that the two quantities describe fundamentally different phenomena. The BCS coherence length represents the average size of a Cooper pair within the superconductor, whereas the effective coherence length of the junction here quantifies the spatial extent of the induced correlations within the normal region.
@@ III.4 Disorder in the SN junction
[…]